OpenStudy (australopithecus):

Find the directional derivative: of f(x,y) = ye^(-x) at point (2,1), theta = pi/4 so I got u = (cos(pi4), sin(pi/4)) f_x(x,y) = -ye^-x f_y(x,y)= e^-x thus I got Duf(2,1) = (-1/e^2)(cos(pi/4)), (1/e^2)(sin(pi/4))) is this correct?

OpenStudy (australopithecus):

looks like i'm wrong :L

OpenStudy (australopithecus):

I'm suppose to add them I think (-1/e^2)(cos(pi/4))+ (1/e^2)(sin(pi/4)))

OpenStudy (australopithecus):

I guess not I get zero if I add them, and the answer is \[2 + \sqrt{3}/2\]

OpenStudy (australopithecus):

oh crud used the wrong coordinates, should be 0,4

OpenStudy (anonymous):

I was already wondering where the mistake could be, what you did is correct isn't it?

OpenStudy (australopithecus):

let me rework out the problem I used the wrong point

OpenStudy (australopithecus):

Wow brain isn't working: Duf(0,4) = -4(cos(pi/4)) + 1(sin(pi/4)))

OpenStudy (australopithecus):

I still don't get the right answer that equals to -3/(2)^(1/2)

OpenStudy (australopithecus):

I used the wrong theta as well lol

OpenStudy (australopithecus):

Duf(0,4) = -4(cos(2pi/3)) + 1(sin(2pi/3)))

OpenStudy (australopithecus):

ok never mind I think I got it

OpenStudy (anonymous):

I remember the good old times where I was 'stuck' at a problem, staring at a solution and unable to derive it, 30-60 minutes later I noticed that I was looking at the solution for a different exercise.

OpenStudy (australopithecus):

lol

OpenStudy (anonymous):

good times, good times. Long evenings (-; well glad you managed it!